# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

### v1

@@ 5.2. Preliminary lemmas
    As in the previous case, we conclude that $\zeta_{U,L}(s)$ has a simple pole at $s=a(L)$.
Since the condition $A_{L}=B_{L}$ (resp. $A_{L}<B_{L}$, $A_{L}>B_{L}$) is equivalent to $\xi=\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$ (resp. $\xi<\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$, $\xi>\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma$), the desired result follows from the properties stated above. This completes the proof of the theorem.
@@ 5.4. Decomposition of Hirzebruch–Kleinschmidt varieties
$\tilde{x}_{ij}:=\left\{\begin{array}[]{ll}x_{ij}&\textup{if }i\leq t^{\prime}\text{ and }j\leq r^{\prime},\\ 0&\textup{otherwise}.\end{array}\right.$
In other words, we enlarge the matrix $(x_{ij})$ of size $t^{\prime}\times r^{\prime}$ to a matrix of size $t\times r$ by adding zero columns on the right and zero rows on the bottom, and similarly we enlarge $[y_{1}:\ldots:y_{t^{\prime}}]$ by adding zeroes on the right.
@@
$\tilde{z}_{ij}:=\left\{\begin{array}[]{ll}z_{j}&\textup{if }i=1,1\leq j\leq r^{\prime}\\ 0&\textup{otherwise},\end{array}\right.$
@@
For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, let us identify $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ with its image under (25). Similarly, let us identify $\mathbb{P}^{r^{\prime}}$ and $\mathbb{P}^{t-1}$ with their images under (26) and (27), which we denote by $\mathbb{P}^{r^{\prime}}_{1}$ and $\mathbb{P}^{t-1}_{2}$, respectively.
@@
Let $X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety. If $r\geq 2$, then we have the disjoint union decomposition
@@
where $P_{0}=([1:(0)_{i\in I_{t},j\in I_{r}}],[1:0:\ldots:0])$. Moreover, given $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big, we have the following properties:
@@
    For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, the restriction of $H_{L}$ to $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ equals $H_{L^{\prime}}$ where $L^{\prime}:=\gamma h^{\prime}+\xi f^{\prime}$ with $\{h^{\prime},f^{\prime}\}$ the basis of $\operatorname{Pic}(X_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}}))$ given in Proposition 3.2.
@@
Assume $r\geq 2$ and let $x=([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])\in X_{d}(a_{1},\ldots,a_{r})$. If $x_{ir}=0$ for all $i\in I_{t}$, then $x\in X_{d-1}(a_{1},\ldots,a_{r-1})$. Assume this is not the case, and let $t^{\prime}:=\max\{i\in I_{t}:y_{i}\neq 0\}$. From equation (8) with $n=t^{\prime}$ and $m=i>t^{\prime}$ we see that
$x_{ij}=0\text{ for all }i\in I_{t}\text{ with }i>t^{\prime}\text{ and all }j\in I_{r}.$
Assume $t^{\prime}=1$. Then, $x_{ij}=0$ for all $i\in I_{t}$ with $i>1$ and all $j\in I_{r}$. Since $x\not\in X_{d-1}(a_{1},\ldots,a_{r-1})$, we have $x_{1r}\neq 0$, and this implies that $x\in\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1}$. Finally, assume $t^{\prime}>1$. Since $y_{t^{\prime}}\neq 0$, having $x_{t^{\prime}r}=0$ would imply $x_{ir}=0$ also for all $i\neq t^{\prime}$ (because of equation (8) with $n=t^{\prime}$ and $j=r$), but this contradicts the fact that $x\not\in X_{d-1}(a_{1},\ldots,a_{r-1})$. So, $x_{t^{\prime}r}\neq 0$ and we have in this case that $x\in U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})$. This proves that
Moreover, these unions are disjoint by construction. This proves (28). The proof of the decomposition (29) in the case $r=1$ is completely analogous. Finally, the statements about the various restrictions of the height function $H_{L}$ follow directly from (11). This proves the lemma. ∎
Remark 5.16.
5.5. The anticanonical height zeta function - Proof of Theorem 1.2

### v2

@@ 5.2. Preliminary lemmas
    As in the previous case, using that $\mathbf{R}_{K}\left(1-t,\left(\gamma a_{r}+\xi\right)A_{L}-((r+1)a_{r}-|\mathbf{a}|)\right)$ is positive, we conclude that $\zeta_{U,L}(s)$ has a pole of order $b(L)=1$ at $s=a(L)$.
This completes the proof of the Theorem 5.2.
5.4. Counting rational points of large height
In this section we use Theorems 5.2 and 5.4 to deduce, via Theorem 2.1, asymptotic formulas for the number of points of large height on a Hirzebruch–Kleinschmidt variety $X$, with respect to height functions $H_{L}$ associated to $L\in\operatorname{Pic}(X)$ big. We start by introducing the following useful definition.
Definition 5.15.
Given a Hirzebruch–Kleinschmidt variety $X$ and $L\in\operatorname{Pic}(X)$ big, we define
$\eta_{L}:=\max\{\eta\in\mathbb{Z}^{+}:\eta^{-1}L\in\operatorname{Pic}(X)\}.$
In other words, $\eta_{L}$ is the unique positive integer such that $\eta_{L}^{-1}L\in\operatorname{Pic}(X)$ and is primitive.
Note that, if we write $L=\gamma h+\xi f$ with $\{h,f\}$ the basis of $\operatorname{Pic}(X)$ given in Proposition 3.2, then $\eta_{L}=\mathrm{g.c.d.}(\gamma,\xi)$. Moreover, it follows from (17) that the height function $H_{L}$ takes values in $q^{\eta_{L}\mathbb{Z}}$, hence the only interesting values of the counting function $\tilde{N}(q^{M},H_{L},X)$ occur when $M$ is divisible by $\eta_{L}$.
In order to state the following theorem, we also define, for $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big, the number
$a^{\prime}(L):=\left\{\begin{array}[]{ll}\max\left\{A_{L}-\frac{1}{\gamma},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},&\text{if }A_{L}=B_{L},\\ \max\left\{A_{L},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},&\text{if }A_{L}<B_{L},\\ \max\left\{A_{L}-\frac{1}{\gamma},B_{L}\right\},&\text{if }A_{L}>B_{L}.\end{array}\right.$
Moreover, recalling that $u$ denotes the unit for the convolution product $\star$ on functions $\operatorname{Div}^{+}(K)\to\mathbb{R}$ (see Definition 5.5), we extend the definitions of $N^{1}_{m}(D)$ and $\tilde{N}_{m}(D)$ given in (23) and (24), in order to include the case $m=0$, by putting $N^{1}_{0}(D):=1$ and $\tilde{N}_{0}(D):=u(D)$. Then, for every integer $m\geq 0$ we have that $(N^{1}_{m},\tilde{N}_{m})$ is a $\mu$-couple and $N_{m}^{1}(D)=q^{m\ell(D)}$. Finally, for every integer $m\geq 0$ we define $F_{m}:\operatorname{Div}^{+}(K)\to\mathbb{R}$ as
$F_{m}(D):=\prod_{\begin{subarray}{c}v\in\operatorname{Val}(K)\\ v(D)>0\end{subarray}}(1-q^{-mf_{v}}).$
In particular, $F_{0}=u$.
The first main theorem of this section gives an asymptotic formula for the number
$\tilde{N}(U,H_{L},q^{M}):=\#\{P\in U(K):H_{L}(P)=q^{M}\},$
for $M\in\eta_{L}\mathbb{Z}$ large, $U$ the good open subset of a Hirzebruch–Kleinschmidt variety $X=X_{d}(a_{1},\ldots,a_{r})$ with $a_{r}>0$ and $L\in\operatorname{Pic}(X)$ big. The case $a_{r}=0$ is treated later in this section.
Theorem 5.16.
Let $X:=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}>0$, $U:=U_{d}(a_{1},\ldots,a_{r})$ the good open subset of $X$ and $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. For every integer $n\in\eta_{L}\mathbb{Z}$ we put $n_{0}:=\frac{n}{\eta_{L}}$. Then, for every $\delta>a^{\prime}(L)$ we have the estimate
$\tilde{N}(U,H_{L},q^{M})=Q_{L}(M)q^{a(L)M}+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\eta_{L}\mathbb{Z},$
where $Q_{L}(M)$ is given as follows:
  1. (1)
    If $A_{L}=B_{L}$, then $Q_{L}(M)=C_{L}M+\tilde{C}_{L}(M)$ with
    $C_{L}=\frac{q^{(d+2)(1-g)}\eta_{L}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)\gamma(q-1)^{2}}>0,$
    and $\tilde{C}_{L}(M)\in\mathbb{R}$ depending only on $L$ and on $M_{0}$ mod $\gamma_{0}(\gamma a_{r}+\xi)_{0}$.
  2. (2)
    If $A_{L}<B_{L}$, then
    $Q_{L}(M)=\frac{q^{(d+2-N_{X})(1-g)}h_{K}}{\zeta_{K}\left(t\right)(q-1)}\sum\limits_{\begin{subarray}{c}D\geq 0\\ \deg(D)\equiv M_{0}\gamma_{0}^{\prime}\,(\gamma a_{r}+\xi)_{0}\end{subarray}}(\tilde{N}_{N_{X}-1}\star F_{r+1-N_{X}})(D)q^{-\deg(D)\left(\gamma B_{L}-r+N_{X}-1\right)},$
    where $\gamma_{0}^{\prime}$ denotes the multiplicative inverse of $\gamma_{0}$ mod $(\gamma a_{r}+\xi)_{0}$ and the sum is over all divisors $D\geq 0$ satisfying $\deg(D)\equiv M_{0}\gamma_{0}^{\prime}\,\text{mod }(\gamma a_{r}+\xi)_{0}$. In particular, $Q_{L}(M)>0$ and it depends only on $L$ and on $M_{0}$ mod $(\gamma a_{r}+\xi)_{0}$.
  3. (3)
    If $A_{L}>B_{L}$, then
    $Q_{L}(M)=\frac{q^{(r+1)(1-g)}h_{K}}{\zeta_{K}(r+1)(q-1)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\equiv M_{0}(\gamma a_{r}+\xi)_{0}^{\prime}\,(\gamma_{0})\end{subarray}}\tilde{N}_{t-1}(D)q^{-\deg(D)((\gamma a_{r}+\xi)A_{L}-((r+1)a_{r}-|\mathbf{a}|))},$
    where $(\gamma a_{r}+\xi)_{0}^{\prime}$ denotes the multiplicative inverse of $(\gamma a_{r}+\xi)_{0}$ mod $\gamma_{0}$ and the sum is over all divisors $D\geq 0$ satisfying $\deg(D)\equiv M_{0}(\gamma a_{r}+\xi)_{0}^{\prime}\,\text{mod }\gamma_{0}$. In particular, $Q_{L}(M)>0$ and it depends only on $L$ and on $M_{0}$ mod $\gamma_{0}$.
The proof of Theorem 5.16 is given below, after the following lemma.
Lemma 5.17.
Let $a,b\in\mathbb{Z}$ with $a\leq 0\leq b$ and $s\in\mathbb{C}$ with $\Re(s)>1+b-a$. Then, the following properties hold:
  1. (1)
    We have
    $\frac{\mathbf{R}_{K}(a,s-b)}{\zeta_{K}(s)}=\sum_{D\geq 0}(\tilde{N}_{|a|}\star F_{b})(D)q^{-(s-b)\deg(D)}.$ (33)
  2. (2)
    If $(a,b)\neq(0,0)$ then for every choice of integers $M,N$ with $N\geq 1$ we have
    $\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\equiv M\,(N)\end{subarray}}(\tilde{N}_{|a|}\star F_{b})(D)q^{-(s-b)\deg(D)}>0,$ (34)
    where the sum is over all divisors $D\geq 0$ satisfying $\deg(D)\equiv M$ mod $N$.
Proof.
By definition of $\mathbf{R}_{K}$ and Corollary 5.9, we have [^6] The hypothesis $\Re(s)>1+b-a$ ensures that all the series appearing in this proof are absolutely convergent.
$\mathbf{R}_{K}(a,s-b)$ $=$ $\sum_{D\geq 0}q^{-a\ell(D)}q^{-(s-b)\deg(D)}$
$=$ $\left(\sum_{D\geq 0}N^{1}_{|a|}(D)q^{-(s-b)\deg(D)}\right)\left(\sum_{D\geq 0}\mu(D)q^{-(s-b)\deg(D)}\right)\zeta_{K}(s-b).$
This implies
$\frac{\mathbf{R}_{K}(a,s-b)}{\zeta_{K}(s)}=\left(\sum_{D\geq 0}\tilde{N}_{|a|}(D)q^{-(s-b)\deg(D)}\right)\frac{\zeta_{K}(s-b)}{\zeta_{K}(s)}.$ (35)
Now, using the Euler product expansion of $\zeta_{K}(s)$, we have
$\frac{\zeta_{K}(s-b)}{\zeta_{K}(s)}=\prod_{v\in\operatorname{Val}(K)}\frac{1-q^{-sf_{v}}}{1-q^{-(s-b)f_{v}}}=\prod_{v\in\operatorname{Val}(K)}\left(1+(1-q^{-bf_{v}})\sum_{n=1}^{\infty}q^{-(s-b)f_{v}n}\right).$
This implies
$\frac{\zeta_{K}(s-b)}{\zeta_{K}(s)}=\sum_{D\geq 0}F_{b}(D)q^{-(s-b)\deg(D)}.$ (36)
[…]
$\sum_{\begin{subarray}{c}n\geq 0\\ n\equiv M(N)\end{subarray}}c_{n}(a,b)q^{-ns},\text{ where }c_{n}(a,b):=q^{nb}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)=n\end{subarray}}(\tilde{N}_{|a|}\star F_{b})(D).$
By definition of $\tilde{N}_{|a|}$ and $F_{b}$, we have $\tilde{N}_{|a|}(D),F_{b}(D)\geq 0$ for all $D\in\operatorname{Pic}^{+}(K)$. This shows that $c_{n}(a,b)\geq 0$. Now, assume $b>0$. Then $q^{nb}F_{b}(D)\geq 1$ for all $D\in\operatorname{Pic}^{+}(K)$ with $\deg(D)\leq n$, hence using that $\tilde{N}_{|a|}(0)\geq 1$ we get
$c_{n}(a,b)\geq q^{nb}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)=n\end{subarray}}F_{b}(D)\geq\#\{D\in\operatorname{Div}^{+}(K):\deg(D)=n\}\geq\#\{v\in\operatorname{Val}(K):f_{v}=n\}.$
Similarly, if $b=0$ and $a<0$ then, recalling that $F_{0}=u$ and $\tilde{N}_{1}=N_{1}^{1}\star\mu$, we get
$c_{n}(a,b)=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)=n\end{subarray}}\tilde{N}_{|a|}(D)\geq\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)=n\end{subarray}}\tilde{N}_{1}(D)\geq\sum_{\begin{subarray}{c}v\in\operatorname{Val}(K)\\ f_{v}=n\end{subarray}}\tilde{N}_{1}(v)\geq\sum_{\begin{subarray}{c}v\in\operatorname{Val}(K)\\ f_{v}=n\end{subarray}}(q^{\ell(v)}-q).$
Then, since the number of valuations $v\in\operatorname{Val}(K)$ with $f_{v}=n$ is positive for large enough $n$ depending only on $K$ (see, e.g., [Ros02, Theorem 5.12]), and $\ell(v)>1$ if $f_{v}>\max\{g,2g-2\}$ by the Riemann–Roch theorem, we conclude that $c_{n}(a,b)>0$ for large enough $n$ provided $(a,b)\neq(0,0)$. This completes the proof of the lemma. ∎
Remark 5.18.
When $(a,b)=(0,0)$ we have $(\tilde{N}_{|a|}\star F_{b})(D)=u(D)$ and (34) holds only when $M\equiv 0$ mod $N$.
Proof of Theorem 5.16.
Put $L_{0}:=\eta_{L}^{-1}L$ and let $M\in\eta_{L}\mathbb{Z}$. Then, we have $\tilde{N}(U,H_{L},q^{M})=\tilde{N}(U,H_{L_{0}},q^{M_{0}})$. Now, applying Theorems 5.2 and 2.1 to $\zeta_{U,H_{L_{0}}}(s)$ we get that, for every $\delta_{0}\in\,]a^{\prime}(L_{0}),a(L_{0})[$ we have the estimate
$\tilde{N}(U,H_{L_{0}},q^{M_{0}})=\log(q)^{b(L_{0})}Q_{0}(M_{0})q^{a(L_{0})M_{0}}+O\left(q^{\delta_{0}M_{0}}\right)\quad\text{as }M_{0}\to\infty,M_{0}\in\mathbb{Z},$
where $Q_{0}(M_{0})$ is as follows:
  1. (1)
    If $A_{L_{0}}=B_{L_{0}}$, then $Q_{0}(M_{0})=C_{1}M_{0}+C_{2}(M_{0})$ with
    $C_{1}=\lim_{s\to 2}(s-2)^{2}\zeta_{U,H_{L_{0}}}(s)=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)_{0}\gamma_{0}(q-1)^{2}\log(q)^{2}},$
    and $C_{2}(M_{0})\in\mathbb{R}$ depending on $M_{0}$ mod $\gamma_{0}(\gamma a_{r}+\xi)_{0}$.
  2. (2)
    If $A_{L_{0}}<B_{L_{0}}$, then
    $Q_{0}(M_{0})=\frac{q^{(d+2-N_{X})(1-g)}h_{K}}{\zeta_{K}\left(t\right)(\gamma a_{r}+\xi)_{0}(q-1)\log(q)}\sum_{k=0}^{(\gamma a_{r}+\xi)_{0}-1}e^{\frac{2\pi iM_{0}k}{(\gamma a_{r}+\xi)_{0}}}\frac{\mathbf{R}_{K}\left(1-N_{X},\gamma_{0}w_{k}-r+N_{X}-1\right)}{\zeta_{K}\left(\gamma_{0}w_{k}\right)}$
    where $w_{k}:=B_{L_{0}}+\frac{2\pi ik}{(\gamma a_{r}+\xi)_{0}\log(q)}$. Moreover, using Lemma 5.17(1) we have
    $\frac{\mathbf{R}_{K}\left(1-N_{X},\gamma_{0}w_{k}-r+N_{X}-1\right)}{\zeta_{K}\left(\gamma_{0}w_{k}\right)}$ $=$ $\sum_{D\geq 0}(\tilde{N}_{N_{X}-1}\star F_{r+1-N_{X}})(D)e^{-\deg(D)\frac{2\pi i\gamma_{0}k}{(\gamma a_{r}+\xi)_{0}}}$
    ×$\hskip 28.45274pt\times q^{-\deg(D)\left(\gamma_{0}B_{L_{0}}-r+N_{X}-1\right)}.$
    This implies
    $Q_{0}(M_{0})=\frac{q^{(d+2-N_{X})(1-g)}h_{K}}{\zeta_{K}\left(t\right)(q-1)\log(q)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\equiv M_{0}\gamma_{0}^{\prime}\,((\gamma a_{r}+\xi)_{0})\end{subarray}}(\tilde{N}_{N_{X}-1}\star F_{r+1-N_{X}})(D)q^{-\deg(D)\left(\gamma_{0}B_{L_{0}}-r+N_{X}-1\right)},$
    and this value is positive by Lemma 5.17(2).
  3. (3)
    If $A_{L_{0}}>B_{L_{0}}$, then
    $Q_{0}(M_{0})=\frac{q^{(r+1)(1-g)}h_{K}}{\zeta_{K}(r+1)\gamma_{0}(q-1)\log(q)}\sum_{k=0}^{\gamma_{0}-1}e^{\frac{2\pi iM_{0}k}{\gamma_{0}}}\frac{\mathbf{R}_{K}\left(1-t,(\gamma a_{r}+\xi)_{0}w_{k}-((r+1)a_{r}-|\mathbf{a}|)\right)}{\zeta_{K}((\gamma a_{r}+\xi)_{0}w_{k}-((r+1)a_{r}-|\mathbf{a}|))}$
    where $w_{k}:=A_{L_{0}}+\frac{2\pi ik}{\gamma_{0}\log(q)}$ in this case. As in the previous case, we use Lemma 5.17(1) to write
    $\frac{\mathbf{R}_{K}\left(1-t,(\gamma a_{r}+\xi)_{0}w_{k}-((r+1)a_{r}-|\mathbf{a}|)\right)}{\zeta_{K}((\gamma a_{r}+\xi)_{0}w_{k}-((r+1)a_{r}-|\mathbf{a}|))}$ $=$ $\sum_{D\geq 0}\tilde{N}_{t-1}(D)e^{-\deg(D)\frac{2\pi i(\gamma a_{r}+\xi)_{0}k}{\gamma_{0}}}$
    ×$\hskip 28.45274pt\times q^{-\deg(D)((\gamma a_{r}+\xi)_{0}A_{L_{0}}-((r+1)a_{r}-|\mathbf{a}|))}$
    and get
    $Q_{0}(M_{0})=\frac{q^{(r+1)(1-g)}h_{K}}{\zeta_{K}(r+1)(q-1)\log(q)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\equiv M_{0}(\gamma a_{r}+\xi)_{0}^{\prime}\,(\gamma_{0})\end{subarray}}\tilde{N}_{t-1}(D)q^{-\deg(D)((\gamma a_{r}+\xi)_{0}A_{L_{0}}-((r+1)a_{r}-|\mathbf{a}|))}.$
    By Lemma 5.17(2) this value is positive.
Then, the desired results follow from these computations and the fact that $A_{L_{0}}=\eta_{L}A_{L},B_{L_{0}}=\eta_{L}B_{L},b(L_{0})=b(L),a(L_{0})=\eta_{L}a(L),a^{\prime}(L_{0})=\eta_{L}a^{\prime}(L)$ and putting $Q_{L}(M):=\log(q)^{b(L)}Q_{0}(M_{0})$. This completes the proof of the theorem. ∎
We now treat the easier case where $a_{r}=0$. In this setting, for $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big, we define
$a^{\prime\prime}(L):=\left\{\begin{array}[]{ll}\max\left\{\frac{1}{2\gamma},\frac{1}{2\xi}\right\},&\text{if }A_{L}=B_{L},\\ \max\left\{A_{L},\frac{1}{2\xi}\right\},&\text{if }A_{L}<B_{L},\\ \max\left\{\frac{1}{2\gamma},B_{L}\right\},&\text{if }A_{L}>B_{L}.\end{array}\right.$
Also, recall that $\tilde{N}(\mathbb{P}^{n},q^{d}):=\{P\in\mathbb{P}^{n}(K):H_{\mathbb{P}^{n}}(P)=q^{d}\}$ (see Section 4).
Theorem 5.19.
Let $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}=0$ and let $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. Then, for every $\delta\in\,]a^{\prime\prime}(L),a(L)[$ we have the estimate
$\tilde{N}(X,H_{L},q^{M})=Q_{L}(M)q^{a(L)M}+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\eta_{L}\mathbb{Z},$
where, using the same notation as in Theorem 5.16, the function $Q_{L}(M)$ is given as follows:
  1. (1)
    If $A_{L}=B_{L}$, then $Q_{L}(M)=C_{L}M+\tilde{C}_{L}(M)$ with
    $C_{L}=\frac{q^{(d+2)(1-g)}\eta_{L}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)\xi\gamma(q-1)^{2}},$
    and $\tilde{C}_{L}(M)\in\mathbb{R}$ depending only on $L$ and on $M_{0}$ mod $\gamma_{0}\xi_{0}$.
  2. (2)
    If $A_{L}<B_{L}$, then
    $Q_{L}(M)=\frac{q^{t(1-g)}h_{K}}{\zeta_{K}\left(t\right)(q-1)}\sum\limits_{\begin{subarray}{c}d\geq 0\\ d\equiv M_{0}\gamma_{0}^{\prime}\,(\xi_{0})\end{subarray}}\tilde{N}(\mathbb{P}^{r},q^{d})q^{-d\gamma B_{L}}.$
    In particular, $Q_{L}(M)>0$ and it depends only on $L$ and on $M_{0}$ mod $\xi_{0}$.
  3. (3)
    If $A_{L}>B_{L}$, then
    $Q_{L}(M)=\frac{q^{(r+1)(1-g)}h_{K}}{\zeta_{K}(r+1)(q-1)}\sum\limits_{\begin{subarray}{c}d\geq 0\\ d\equiv M_{0}\xi_{0}^{\prime}\,(\gamma_{0})\end{subarray}}\tilde{N}(\mathbb{P}^{t-1},q^{d})q^{-d\xi A_{L}}.$
    In particular, $Q_{L}(M)>0$ and it depends only on $L$ and on $M_{0}$ mod $\gamma_{0}$.
Proof.
The result follows from Theorems 5.4 and 2.1, and is analogous to the proof of Theorem 5.16. Hence, we omit the details for brevity. Here, we only mention that, in the case $A_{L}\neq B_{L}$, the non-vanishing of $Q_{L}(M)$ follows from the fact that $\tilde{N}(\mathbb{P}^{n},q^{d})>0$ for large $d\in\mathbb{Z}$ depending only on $n$ and $K$, because of (16). This proves the theorem. ∎
@@ 5.5. Decomposition of Hirzebruch–Kleinschmidt varieties
$\tilde{x}_{ij}:=\left\{\begin{array}[]{ll}x_{1j}&\textup{if }j\leq r^{\prime},a_{j}=0,i\leq t,\\ x_{ij}&\textup{if }j\leq r^{\prime},a_{j}>0,i\leq r^{\prime},\\ 0&\textup{otherwise}.\end{array}\right.$
@@
$\tilde{z}_{ij}:=\left\{\begin{array}[]{ll}z_{j}&\textup{if }j\leq r^{\prime},a_{j}=0,i\leq r,\textup{ or }j\leq r^{\prime},a_{j}>0,i=1,\\ 0&\textup{otherwise},\end{array}\right.$
@@
For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, let us identify $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ and $X_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ with their images under (37). Similarly, let us identify $\mathbb{P}^{r^{\prime}}$ and $\mathbb{P}^{t-1}$ with their images under (38) and (39), which we denote by $\mathbb{P}^{r^{\prime}}_{1}$ and $\mathbb{P}^{t-1}_{2}$, respectively.
@@
Let $X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety with $a_{r}>0$. If $r\geq 2$, then we have the disjoint union decomposition
@@
where $P_{0}=([1:(0)_{i\in I_{t},j\in I_{r}}],[1:0:\ldots:0])$.
Proof.
Assume $r\geq 2$ and let $x=([x_{0}:(x_{ij})_{i\in I_{t},j\in I_{r}}],[y_{1}:\ldots:y_{t}])\in X_{d}(a_{1},\ldots,a_{r})$. If $x_{ir}=0$ for all $i\in I_{t}$, then $x\in X_{d-1}(a_{1},\ldots,a_{r-1})$. Assume this is not the case, and let $t^{\prime}:=\max\{i\in I_{t}:x_{ir}\neq 0\}$. From equation (14) with $j=r$ and $m=t^{\prime}$ we see that $y_{n}=0$ for all $n>t^{\prime}$ (because $a_{r}>0$). At the same time we see that having $y_{t^{\prime}}=0$ would imply $y_{n}=0$ also for all $n\neq t^{\prime}$, which is impossible since $[y_{1}:\ldots:y_{t}]\in\mathbb{P}^{t-1}$. Hence, $y_{t^{\prime}}\neq 0$. Now, from equation (14) with $n=t^{\prime}$ and $m=i>t^{\prime}$ we see that
$x_{ij}=0\text{ for all }i\in I_{t}\text{ with }i>t^{\prime}\text{ and all }j\in I_{r}\text{ with }a_{j}>0.$
If $t^{\prime}\geq 2$, then this shows that $x\in U_{t^{\prime}+r-1}(a_{1},\ldots,a_{r})$. If $t^{\prime}=1$, then we conclude that $x\in\mathbb{P}^{r}_{1}\setminus\mathbb{P}^{r-1}_{1}$. This proves that
Moreover, these unions are disjoint by construction. This proves (40). The proof of the decomposition (41) in the case $r=1$ is completely analogous. This proves the lemma. ∎
Remark 5.21.
In the case $a_{r}=0$ we have the following simpler decomposition.
Lemma 5.22.
Let $X_{d}(a_{1},\ldots,a_{r})\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ be a Hirzebruch–Kleinschmidt variety with $a_{r}=0$. If $r\geq 2$, then we have the disjoint union decomposition
$X_{d}(a_{1},\ldots,a_{r})=X_{d-1}(a_{1},\ldots,a_{r-1})\sqcup U_{d}(a_{1},\ldots,a_{r}).$
If $r=1$, then we have the disjoint union decomposition
$X_{d}(a_{1})=\mathbb{P}^{t-1}_{2}\sqcup U_{d}(a_{1}).$
Proof.
Write $x\in X_{d}(a_{1},\ldots,a_{r})\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ as $x=([x_{0}:x_{1}:\ldots:x_{r}],[y_{1},\ldots,y_{t}])$. Then, the components in the above decompositions correspond to the cases $x_{r}=0$ and $x_{r}\neq 0$. This proves the lemma. ∎
Finally, we state the following lemma regarding the restriction of a height function on a Hirzebruch–Kleinschmidt variety, to each component in the above decompositions.
Lemma 5.23.
Given $L=\gamma h+\xi f\in\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ big, we have the following properties:
@@
    For $1\leq r^{\prime}\leq r$ and $2\leq t^{\prime}\leq t$, the restriction of $H_{L}$ to $U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ corresponds to $H_{L^{\prime}}$ where $L^{\prime}:=\gamma h^{\prime}+\xi f^{\prime}$ with $\{h^{\prime},f^{\prime}\}$ the basis of $\operatorname{Pic}(X_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}}))$ given in Proposition 3.2.
@@
These statements are a direct consequence of (17). ∎
5.6. The anticanonical height zeta function - Proof of Theorem 1.4 and Corollary 1.6